### v2

Abstract
The backpropagation algorithm remains the dominant and most successful method for training deep neural networks (DNNs). At the same time, training DNNs at scale comes at a significant computational cost and therefore a high carbon footprint. Converging evidence suggests that input decorrelation may speed up deep learning. However, to date, this has not yet translated into substantial improvements in training efficiency in large-scale DNNs. This is mainly caused by the challenge of enforcing fast and stable network-wide decorrelation. Here, we show for the first time that much more efficient training of very deep neural networks using decorrelated backpropagation is feasible. To achieve this goal we made use of a novel algorithm which induces network-wide input decorrelation using minimal computational overhead. By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training a 18-layer deep residual network. This demonstrates that decorrelation provides exciting prospects for efficient deep learning at scale.
keywords:
Deep learning, Backpropagation, Efficient AI, Decorrelation, Whitening
1 Introduction
Modern AI relies heavily on deep learning (DL), which refers to the training of very deep neural network (DNN) models using massive datasets deployed on high-performance compute clusters [1]. The established way of implementing learning in artificial neural networks is through the backpropagation (BP) algorithm [2, 3]. BP is a gradient-based method that implements reverse-mode automatic differentiation to compute the gradients needed for parameter updating in neural networks [4].
At the same time, training of DNNs consisting of many layers of nonlinear transformations is computationally expensive and has been associated with significant energy consumption at a global scale [5, 6, 7, 8, 9, 10]. Hence, we are in need of green AI solutions which can significantly reduce the energy consumption of modern AI systems [11, 12, 13]. Patterson et al. [14] describe several ways to reduce the carbon consumption of deep learning, focusing mostly on more effective deployment methods. Another key reason for deep learning’s large carbon consumption, however, is the inefficiency of the backpropagation algorithm itself. BP requires many iterations of gradient descent steps until parameters converge to their optimal values. For example, training a large GPT model can easily take several weeks on a large compute cluster [15].
In this paper, we show that deep learning can be made much more efficient by enforcing decorrelated inputs throughout the network. Decorrelation has previously been proposed to make credit assignment more efficient in neural networks [16]. Intuitively, if a network layer’s inputs have highly correlated features, it will be more difficult for the learning algorithm to perform credit assignment as it is now unclear if a change in the loss should be attributed to feature $i$ or feature $j$ in case both are correlated. Previous work has indeed shown that promoting decorrelation positively impacts training in relatively shallow networks [17].
Similar results have been obtained when inputs are whitened such that inputs are not only decorrelated but also forced to have unit variance [18]. In [19], whitening has been shown to have a positive impact in training deep networks, resulting in faster convergence when measured by number of epochs and somewhat lower test error compared to regular BP. However, due to computational overhead this did not yet translate into a substantial reduction in wall-clock training time. Desjardins et al. [20] also demonstrate that whitening has benefits when training DNNs, though this did not result in faster convergence to the minimal validation loss compared to standard BP using batch normalisation. At the same time, [21] suggest that strict whitening can have a negative impact on generalization performance.
Hence, while the potential benefits of decorrelation and whitening are encouraging, their application for significantly reducing real-world training times in modern deep neural networks remains to be shown. This has two main reasons. First, enforcing network decorrelation during training is associated with significant computational overhead, thereby eliminating any gains in training efficiency. Second, it is not obvious how to effectively implement decorrelation in a stable manner in deep (convolutional) neural networks.
In this paper we demonstrate, for the first time, that training of large-scale DNNs can be made much more efficient through the development of a novel decorrelated backpropagation (DBP) algorithm. DBP combines automatic differentiation with an efficient iterative local learning rule which effectively decorrelates layer inputs across the network. This iterative decorrelation procedure was previously introduced by Ahmad et al. [17] and has been shown to speed up learning in limited-depth fully-connected neural networks. Here, we extend the procedure to more efficiently and effectively train deep residual networks [22]. This is achieved by making the procedure suitable for application in convolutional layers and ensuring that decorrelation remains stable across layers. Furthermore, we show how the same algorithm can be used to whiten layer inputs, allowing us to compare the efficacy of decorrelation and whitening in DNNs.
In the following, we show that DBP yields a two-fold reduction in training time, while achieving better performance compared to regular BP. Hence, widespread application of our approach can yield a substantial reduction in the carbon consumption of modern deep learning.
2 Methods
2.1 Decorrelated backpropagation
Let us consider a DNN consisting of $K$ layers implementing some parameterized nonlinear transform. Deep learning typically proceeds via reverse-mode automatic differentiation, also known as the backpropagation (BP) algorithm. Backpropagation updates the parameters $\bm{\theta}$ (weights and biases) of a network according to
$\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}\,,$ (1)
where $\eta$ is the learning rate, $\bm{\theta}$ are the network parameters and $\nabla_{\bm{\theta}}\mathcal{L}$ is the gradient of the loss.
Decorrelated backpropagation (DBP) proceeds as regular backpropagation but additionally enforces decorrelated inputs to layers in the network. Since decorrelation is imposed independently in each layer, we concentrate on describing the decorrelation procedure for one such layer, implementing a nonlinear transformation
$\mathbf{y}=f\left(\mathbf{W}\mathbf{x}\right)$ (2)
with $\mathbf{W}\in\mathbb{R}^{d\times d}$ the weights and $\mathbf{x}\in\mathbb{R}^{d}$ the input. This transformation may be either the kernel function of a convolutional layer applied to a (flattened) input patch or the transformation in a fully-connected layer. Note further that we ignore biases without loss of generality since these can be represented using weights with fixed constant input. Input decorrelation refers to the property that $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is diagonal for inputs $\mathbf{x}$, where the expectation is taken over the data distribution. In case of whitened inputs, $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is assumed to be the identity matrix [23]. We will refer to the former as the covariance constraint and the latter as the variance constraint.
To ensure that the input $\mathbf{x}$ is decorrelated, we assume that it is the result of a linear transform
$\mathbf{x}=\mathbf{R}\mathbf{z}\,,$ (3)
where the decorrelating matrix $\mathbf{R}$ transforms the correlated input $\mathbf{z}$ into a decorrelated input $\mathbf{x}$. Hence, the transformation in a layer is described by
$\mathbf{y}=f(\mathbf{W}\mathbf{R}\mathbf{z})=f(\mathbf{W}\mathbf{x})=f(\mathbf{A}\mathbf{z})$ (4)
with $\mathbf{A}=\mathbf{W}\mathbf{R}$. The output $\mathbf{y}_{l}$ of the $l$-th layer becomes the correlated input $\mathbf{z}_{l+1}$ to the next layer.
2.2 Decorrelation learning rule
To implement decorrelated backpropagation, we need to update individual decorrelation matrices for each of the layers in the network. This can be achieved by minimizing $K$ decorrelation loss functions in parallel to minimizing the BP loss with the aim of accelerating learning speed.
One way to achieve whitening as a restricted form of decorrelation is via established approaches such as zero-phase component analysis (ZCA) [24, 25]. However, naive application of ZCA is prohibitively costly in practice since it requires application of the algorithm to the inputs of each layer for all of the examples in the dataset at each step of gradient descent. More efficient batch implementations of whitening have been shown to have a positive impact on neural network training [18, 20, 19] but this did not yet yield a substantial reduction in training time in large-scale models when considering the moment at which minimal validation loss is achieved. See Appendix A for a complexity analysis of the different approaches.
To enable efficient decorrelation and whitening, we extend a recently introduced approach that enables more efficient network-wide decorrelation [17] and make this suitable for application in large-scale models. To derive our decorrelation learning rule, we start by defining the total decorrelation loss $\langle L\rangle$ as the average decorrelation loss across layers. The layer-specific decorrelation loss is defined as $L=\langle\ell(\mathbf{x})\rangle$, where the expectation is taken over the empirical distribution and $\ell(\mathbf{x})=\sum_{i=1}^{d}\ell_{i}(\mathbf{x})$ is the sum over unit-wise losses $\ell_{i}$. We define the unit-wise loss as
$\ell_{i}(\mathbf{x})=(1-\kappa)\frac{1}{2}\sum_{j\neq i}(x_{i}x_{j})^{2}+\kappa\frac{1}{4}(x_{i}^{2}-1)^{2}\,,$ (5)
where the regularization parameter $\kappa$ interpolates between the covariance constraint ($\kappa=0$) and the unit variance constraint ($\kappa=1$) such that that $\kappa\in(0,1)$ promotes whitening. It follows that we may write the layer-wise decorrelation loss compactly as
$\ell(x)=\left\langle(1-\kappa)\textrm{Tr}(\mathbf{C}\mathbf{C}^{\top})+\kappa\textrm{Tr}(\mathbf{V}\mathbf{V}^{\top})\right\rangle\,,$ (6)
where $\mathbf{C}=\mathbf{x}\mathbf{x}^{\top}-\mathbf{D}$ with $\mathbf{D}=\textrm{diag}(x_{1}^{2},\ldots,x_{d}^{2})$ and $\mathbf{V}=\textrm{diag}(x_{1}^{2}-1,\ldots,x_{d}^{2}-1)$.
Since the decorrelation loss is minimized independently in each layer, we concentrate on describing the minimization of the decorrelation loss for one such layer, as in Eq. 2. Let us first consider decreasing the correlation between variables for one input vector $\mathbf{x}$. This can be achieved using an update step of the form
$\mathbf{x}\leftarrow\mathbf{x}-\epsilon\nabla_{\mathbf{x}}\ell$ (7)
with $\epsilon$ the decorrelation learning rate. To compute the gradient, we first consider the partial derivative of the unit-wise loss with respect to the input $x_{i}$ given by
$\frac{\partial\ell_{i}}{\partial x_{i}}=(1-\kappa)\sum_{j\colon j\neq i}\left(x_{i}x_{j}\right)x_{j}+\kappa(x_{i}^{2}-1)x_{i}\,.$ (8)
If we vectorize across units, we obtain $\nabla_{\mathbf{x}}\ell=(1-\kappa)\mathbf{C}\mathbf{x}+\kappa\mathbf{V}\mathbf{x}$. If we plug this into the update step in Eq. 7, we obtain
$\mathbf{x}\leftarrow\mathbf{x}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{x}\,.$ (9)
Ultimately, we aim to derive an update rule for $\mathbf{R}$ instead of $\mathbf{x}$. This can be obtained using the identity $\mathbf{x}=\mathbf{R}\mathbf{z}$ since this allows us to write $\mathbf{R}\mathbf{z}\leftarrow\mathbf{R}\mathbf{z}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{R}\mathbf{z}$. Dividing by $\mathbf{z}$ on both sides and averaging over input samples we obtain our decorrelation learning rule
$\mathbf{R}\leftarrow\mathbf{R}-\epsilon\left\langle(1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right\rangle\mathbf{R}\,.$ (10)
This decorrelation rule allows for efficient batch updating of $\mathbf{R}$ since it only requires computing the decorrelated input covariance followed by multiplication with the decorrelation matrix. Note that for $\kappa=0$ we only decorrelate the input and for $\kappa=1$ we only enforce the unit variance constraint. For $10<\kappa<$ 1 we include both constraints to induce input whitening. Figure 1 shows effective decorrelation and whitening on two-dimensional input data.
Figure 1: Demonstration of the decorrelation rule on correlated input data consisting of 1000 examples and two covariates with decorrelation learning rate $\epsilon=0.001$. a) Decorrelation using $\kappa=0$. b) Whitening using $\kappa=0.5$. Mean variance and covariance reported for different iterations.
2.3 Application to deep residual networks
Effective training of deep neural networks using DBP requires several modifications that improve algorithm stability, learning speed and allow application to convolutional rather than fully connected layers. As a canonical example, we consider an 18-layer deep residual network (ResNet). A ResNet consists of multiple residual blocks that implement modern network components such as convolutional layers and residual connections [22]. In the following, we describe the modifications that are required for effectively applying DBP in deep neural networks.
Figure 2: Implementation of decorrelated backpropagation in residual networks. a) Residual blocks as implemented by our networks. b) Patch-wise flattening of the inputs with a flattened dimension of $d=M\times M\times C_{in}$. c) Decorrelating/whitening transform of the data by decorrelation matrix $\mathbf{R}$. d) 1 $\times$ 1 convolution operation with the weights $\mathbf{W}$ on the decorrelated input patches.
First, we apply the decorrelation learning rule to convolutional layers as follows. Consider a convolutional layer with input size $S=M_{in}\times M_{in}\times C_{in}$, where $M_{in}\times M_{in}$ is the size of the feature map and $C_{in}$ is the channel dimension. Naive implementation of decorrelation would require updating a $S\times S$ decorrelation matrix $\mathbf{R}$, which is computationally too prohibitive in case of large feature maps. Instead, as shown in Fig. 2, we do not decorrelate the layer’s entire input, but only the local image patches, so that for an image patch with dimensionality $D=K\times K\times C_{in}$, matrix $\mathbf{R}$ would be of much smaller size $D\times D$. The output of this patchwise decorrelation operation would then be $D$ for every image patch, after which we can apply a $1\times 1$ convolutional kernel of size $D\times C_{out}$ for the forward pass. This local decorrelation is sufficient to ensure more efficient learning of the kernel weights. Figure 2 depicts the structure of a residual block that is extended to implement decorrelating transforms.
Second, applying the decorrelation learning rule can become costly when averaging over input samples in Eq. 10 in case of convolutional layers. This is due to the need to compute an outer product $\mathbf{X}\mathbf{X}^{\top}$ with $\mathbf{X}\in\mathbb{R}^{D\times p}$ the input to a layer, where $p$ is the product of the number of batch elements times the number of patches. We found empirically that updates of $\mathbf{R}$ need not use the entire mini-batch to learn the correlational structure of the data. Sampling just 10% of the samples in each batch yields almost identical performance while significantly reducing the computational overhead of computing $\mathbf{R}$’s updates.
Third, another optimization that decreases the computational overhead of the decorrelating transform is to combine the matrices $\mathbf{W}$ and $\mathbf{R}$ into a condensed matrix $\mathbf{A}=\mathbf{W}\mathbf{R}$ prior to multiplying by the correlated input $\mathbf{z}$. The dimensionality of $\mathbf{z}$ is much larger than that of $\mathbf{W}$ due to the batch dimension, and thus we replace an expensive multiplication $\mathbf{W}(\mathbf{R}\mathbf{z})$ by a cheaper multiplication $(\mathbf{W}\mathbf{R})\mathbf{z}$, significantly reducing the time required for performing a forward pass. As an additional benefit, when training is complete, we only need to store the $\mathbf{A}$ matrices.
2.4 Experimental validation
To evaluate learning performance, an 18-layer deep ResNet model was trained and tested on the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) dataset [26]. This dataset spans 1000 object classes and contains 1,281,167 training images, 50,000 validation images and 100,000 test images [26].
Data preprocessing consisted of the following steps. First, images were normalized by subtracting the means from the RGB-channels’ values and dividing by their standard deviation. Next, images were rescaled to $256\times 256$ and a $224\times 224$ crop was taken from the center. We used no data augmentation, but shuffled the order of the data every epoch.
All models and algorithms were implemented in PyTorch [27] and run on a compute cluster using Nvidia A100 GPUs and Intel Xeon Platinum 8360Y processors. To initialize the weights of our models, we set $\mathbf{R}$ to the identity matrix and used He initialization [28] for $\mathbf{W}$. Models were trained for 50 epochs to minimize the categorical cross-entropy loss. Exploratory analysis revealed negligible performance difference as a function of batch size and a batch size of 256 was chosen. Instead of using the full batch of images in each decorrelation update step, we reduce the number of samples to 10%. We empirically found this value to show negligible loss in performance, while significantly speeding up runtime. The accuracy reported is the top-1 performance of the models. Reported wall-clock time was measured as the runtime of the training process, excluding performance testing.
Performance of DBP using either decorrelation ($\kappa=0$) or whitening ($\kappa=0.5$) was compared against regular backpropagation. For updating $\mathbf{W}$, we use the Adam optimizer [29] with a learning rate of $\eta=1.6\cdot 10^{-4}$ and optimizer parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We added a small value of $10^{-8}$ to the denominator of the updates for numerical stability. The decorrelation matrix $\mathbf{R}$ is updated with stochastic gradient descent using a learning rate of $\epsilon=10^{-5}$. To ensure a fair comparison between algorithms, we made sure that all algorithms attained their highest convergence speed by performing a two-dimensional grid search over the BP ($\eta$) and decorrelation ($\epsilon$) learning rates. See Section B for the grid-search results.
3 Results
In the following, we analyse the performance of BP versus DBP on the 18-layer ResNet model trained on the ImageNet task.
3.1 DBP effectively decorrelates inputs to all network layers
In Fig. 3, the change is input correlation is shown for each layer during the training process for DBP. To quantify input correlation we use the mean squared values of the strictly lower triangular part of the $\mathbf{C}$ matrix. It can be seen that layer inputs become rapidly decorrelated during the first 10 epochs and correlations remain consistently low. This demonstrates that the decorrelating learning rule indeed effectively learns to decorrelate the inputs to all network layers.
Figure 3: Input decorrelation when training a ResNet-18 model on ImageNet for 50 epochs. Network layers are ordered from left to right and from top to bottom. Panel titles indicate the layer type.
3.2 DBP converges much faster than BP
We observe a significant improvement when training the ResNet model using DBP compared to regular BP. As shown in Fig. 4, DBP’s train accuracy (Fig. 4a) improves much faster than that of BP, indicating more effective training (computational work per epoch). The same observation holds for test accuracy (Fig. 4b). Additionally, DBP’s test accuracy (55.2%) also peaks above BP’s test accuracy (54.1%), while doing so in roughly half the number of epochs. Finally, we observe that a slightly higher test accuracy is achieved for whitening ($\kappa=0.5$) compared to decorrelation ($\kappa=0$), demonstrating the added benefit of whitening for this task.
The same patterns can be observed for the loss curves in Appendix C. Furthermore, as shown in Appendix D, a slightly lower test loss at equal train loss is obtained for DBP compared to BP, which shows that this does not come at the expense of generalization performance.
Figure 4: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening). Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars. a) Train accuracy as a function of the number of epochs. b) Test accuracy as a function of the number of epochs.
3.3 DBP training can be achieved at shorter wall-clock times than BP
Recall that DBP requires additional computation in each epoch due to application of the decorrelation update rule in Eq. 10. This implies that faster convergence as a function of the number of epochs may not necessarily translate into more efficient training. Therefore, we also compare wall-clock times in Fig. 5.
Figure 5: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening). Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars. a) Train accuracy as a function of wall-clock time. b) Test accuracy as a function of wall-clock time.
[…]
3.4 DBP significantly reduces carbon emission
Figure 6 visualizes what the efficiency gains are when using DBP ($\kappa=0.5$) over BP in terms of carbon emission.
Figure 6: Efficiency gains when using DBP over BP. a) Training time required relative to BP’s time to maximal accuracy to reach a specific percentage of the maximum performance of DBP. b) Estimated reduction in carbon consumption when using DBP over BP. Estimate produced using https://github.com/mlco2/codecarbon. Dashed line indicates the moment when performance is equal to maximal BP test accuracy.
[…]
Figure 6b provides a measurement of the associated reduction in carbon emission if we run the model to achieve a specific percentage of the DBP test accuracy. The figure shows that a reduction of 640 grams of carbon emission is achieved if we train DBP until it reaches 100% of the BP test accuracy. If we run DBP until it reaches 100% of the DBP test accuracy, we still achieve a reduction of 418 grams of carbon emission.
4 Discussion
Increasing the efficiency of deep learning is of key importance if we are to reduce the carbon footprint of AI. In this paper, we have shown that decorrelated backpropagation provides a viable path towards more efficient deep learning. Results show that, by replacing BP with DBP, we can achieve higher test accuracy using about half the number of training epochs. Results also show that both decorrelation ($\kappa=0$) and whitening ($\kappa=0.5$) provide similar performance gains, with whitening yielding slightly better results. This does not come at the expense of generalization performance, as shown in Fig. 9. We also demonstrated a substantial reduction in carbon emission when using DBP over BP when training deep neural networks. Global carbon emissions can therefore be substantially reduced when applying DBP at scale.
These results indicate that performance is maximized for some optimal value of $\kappa$ which most effectively balances the covariance and variance constraints. This optimal value is task-dependent. Appendix F shows that training a medium-sized neural network on the CIFAR10 task yields similar performance gains but now the optimal performance is obtained for $\kappa=0$.
It should be mentioned that network-wide decorrelation comes with a number of additional advantages, which we did not further explore here. As shown in [17], decorrelated inputs allow for rapid linear approximation of non-linear functions, allowing for filter visualisation in deep networks with applications in explainable AI [30, 31]. It also allows for network compression at virtually no computational overhead, which may further decrease carbon consumption at inference time [13]. Intriguingly, input decorrelation and whitening have also been shown to be a feature of neural processing [32, 25, 33, 34, 35, 36, 37]. The work presented here suggests that decorrelation may be an important factor when considering synaptic plasticity mechanisms.
A number of considerations do need to be taken into account when using DBP. First, DBP requires updating and storing of decorrelation matrices needed for input decorrelation. Hence, the algorithm introduces some computational overhead and increased memory requirements. This is reflected by the increased wall-clock time per epoch. To mitigate this, more effective use could be made of the sparseness structure of the decorrelation matrix. For instance, rather than learning a full matrix $\mathbf{R}$ we may choose to learn a lower-triangular matrix, which can be achieved by using the strictly lower triangular part of $\mathbf{C}$ in Eq. 10. Other approaches to reduce wall-clock time may be the use of low-rank approximations of $\mathbf{R}$, using higher layer-specific decorrelation learning rates, or only incorporating decorrelation in that subset of the layers which have the biggest impact in terms of reducing convergence time [19].
Second, DBP needs proper fine-tuning since the stability of the algorithm depends on the two learning rates $\eta$ and $\epsilon$, whose ratio needs to be set appropriately such that the decorrelation rate throughout all network layers remains optimal. We empirically observe that this fine-tuning depends on the employed dataset, network architecture and loss function. Additional work on the interaction between these factors may provide insights into how to optimally set these parameters.
An important consideration here is scaling the updates according to layer size. Note that Eq. 10 ignores the layer size in the normalisation, which implies that decorrelation updates have a larger impact in larger layers whereas whitening induced by $\kappa>0$ has a smaller impact in larger layers. In simulation work, we observe that such normalisation ensures that the impact of decorrelation updates is comparable for different layer sizes. However, in our empirical work using very deep neural networks, we find that ignoring this normalisation actually improves performance. There are different reasons why this may be the case. First, the resulting stronger decorrelation in deep large layers when ignoring normalisation may have a beneficial impact since deeper layers are more strongly affected by multiple preceding nonlinearities. Second, the decorrelation loss that is minimized by Eq. 10 may not be the optimal metric to ensure fastest convergence. That is, there may be a discrepancy between the decorrelation gradient direction and the optimal gradient direction to align the inputs.
In this work, we demonstrated a significant speedup in deep residual networks, which are state-of-the-art models for computer vision. In our follow-up work, we aim to explore the efficiency gains that can be obtained for other computational tasks and network architectures. Exciting prospects are the use of decorrelated backpropagation for deep reinforcement learning [38, 39] or training of foundation models [40], both of which are notoriously resource intensive. We further expect that additional theoretical work on the optimal alignment of layer inputs combined with efficient low-level implementations will yield even larger gains in convergence speed. Concluding, by demonstrating the efficiency of decorrelated backpropagation in modern deep neural networks, we hope to contribute to reducing the energy consumption of modern AI systems.
Code availability
Code for training and evaluating the decorrelated backpropagation algorithm can be downloaded from https://github.com/artcogsys/decorbp.
Acknowledgements
This publication was partly financed by the Dutch Research Council (NWO) via the Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 and a TTW grant with project number 2022/TTW/01389357. We thank Mustafa Acikyurek for his support, SURF (www.surf.nl) for providing access to the National Supercomputer Snellius and Hewlett Packard for providing a Z8 workstation to support this work.
Appendix A Comparison with zero-phase component analysis
Zero-phase component analysis (ZCA) [24, 25] is defined by the whitening transform $\mathbf{x}=\mathbf{C}\mathbf{x}_{c}$, where $\mathbf{x}_{c}=\bar{\mathbf{x}}-\bm{\mu}$ with $\bm{\mu}$ the mean over datapoints and $\mathbf{C}$ is the whitening transform. Let $\mathbf{X}$ be the $n\times d$ matrix consisting of $n$ $d$-dimensional centered inputs. ZCA uses
$\mathbf{C}=\mathbf{\Lambda}^{-1/2}\mathbf{U}^{\top}$ (11)
with $\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$ the eigendecomposition of the covariance matrix $\mathbf{\Sigma}=\mathbf{X}\mathbf{X}^{\top}/n$.
Desjardins et al. [20] propose to periodically (during training) measure the correlation, $\mathbf{C}$, at every layer of a deep neural network and to then apply eigendecomposition method to compute the ZCA transform. This approach to achieving an exact whitening transform is an alternative to what we attempt in this work but has a number of drawbacks. First, the complexity of computing the eigendecomposition itself has a cost equivalent to a matrix multiplication - meaning that this step alone is equally expensive to our decorrelation method. Second, atop this decomposition, Desjardins et al. [20] further compute a matrix inverse of this transform which allows them to periodically compute an exact ZCA-based decorrelation within each layer of their network while also undoing its impact on the layer’s computation (by modifying the layer’s weight matrix $\mathbf{W}$) to ensure that the newly introduced transform does not change the network’s output. We avoid the second step by simultaneously optimising for both the decorrelation and task at every layer and at every update step, without directly computing the ZCA transform. This provides the benefit of not requiring any matrix inversion computation and also ensures that we are continuously optimising the decorrelation transform rather than doing so at fixed intervals.
Appendix B Hyper-parameter optimization
A grid search across learning rate hyper-parameters was performed to ensure that each algorithm reached its optimal performance. An exploratory analysis revealed that although $3.2\cdot 10^{-4}$ gave slightly better performance after five epochs for BP and DBP, later in training a learning rate of $1.6\cdot 10^{-4}$ was better. Therefore, the latter was chosen for all algorithms, which had the convenient side effect of using the same learning rate for all algorithms, making comparison more straightforward. Figure 7 depicts the grid search outcomes.
Figure 7: Train accuracy after five epochs for DBP when varying the BP learning rate $\eta$ and decorrelation learning rate $\epsilon$. Note that a decorrelation learning rate of zero corresponds to regular BP. a) Results using decorrelation ($\kappa=0$). b) Results using whitening ($\kappa=0.5$).
Appendix C Loss curves
Figure 8 reports the loss curves as a function of epochs and wall-clock time for the different algorithms, corresponding to the accuracies reported in the main text.
Figure 8: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening). Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars. a) Train loss as a function of the number of epochs. b) Test loss as a function of the number of epochs. c) Train loss as a function of wall-clock time. d) Test loss as a function of wall-clock time.
Appendix D Train versus test loss comparison
Figure 9 shows the train versus test losses for a ResNet-18 model trained on ImageNet for both BP and DBP. As DBP’s test loss is slightly lower than the BP test loss at the same value of the train loss, we conclude that DBP yields slightly better generalization performance compared to BP.
Figure 9: Scatterplot of train versus test loss for BP and DBP.
Appendix E Peak performance comparison
Table 1 provides an overview of peak performance for the different algorithms, as well as the epoch number and time at which peak test performance was achieved.
Table 1: Peak performance for BP, DBP in terms of loss and accuracy. Epoch and Time indicate the epoch number and wall-clock time at which peak accuracy was achieved. Best results shown in boldface, where we focus on test results as this quantifies generalization performance.
Loss Accuracy Epoch Time (hours)
Method $\kappa$ Train Test Train Test Train Test Train Test
BP - 1.24 2.90 99.4 54.1 49 12 16.0 3.7
DBP 0 1.21 2.86 99.5 54.8 50 7 18.4 2.3
DBP 0.5 1.22 2.84 99.5 55.2 50 7 18.4 2.3
Appendix F Impact of $\kappa$ parameter on CIFAR10 performance
Figure 10 shows train and test performance for a small three-layer ConvNet trained on CIFAR10 for BP and for several settings of DBP where the $\kappa$ parameter was varied. Setting $\kappa=0$ reached the highest peak test accuracy for this particular task. Performance gains for this smaller model are even more pronounced yet result in a smaller reduction in carbon consumption due to the model’s smaller size. Note further that peak test performance for DBP is larger than that of BP, again indicating improved generalization performance. A batch size of 256 and learning rate of $1\times 10^{-3}$ were used in all experiments. The decorrelation/whitening learning rate was also set to $1\times 10^{-3}$.
Figure 10: Performance on CIFAR10 for BP and several $\kappa$ settings of DBP. a) Train accuracy. b) Test accuracy.

### v4

@@
-
+Efficient Deep Learning with Decorrelated Backpropagation
 Abstract
@@
 This is mainly caused by the challenge of enforcing fast and stable network-wide decorrelation.
-Here, we show for the first time that much more efficient training of very deep neural networks using decorrelated backpropagation is feasible.
+Here, we show for the first time that much more efficient training of deep convolutional neural networks is feasible by embracing decorrelated backpropagation as a mechanism for learning.
 To achieve this goal we made use of a novel algorithm which induces network-wide input decorrelation using minimal computational overhead.
-By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training a 18-layer deep residual network.
+By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training several deep networks up to a 50-layer ResNet model.
 This demonstrates that decorrelation provides exciting prospects for efficient deep learning at scale.
-keywords:
-Deep learning, Backpropagation, Efficient AI, Decorrelation, Whitening
 1 Introduction
-Modern AI relies heavily on deep learning (DL), which refers to the training of very deep neural network (DNN) models using massive datasets deployed on high-performance compute clusters [1].
-The established way of implementing learning in artificial neural networks is through the backpropagation (BP) algorithm [2, 3].
-BP is a gradient-based method that implements reverse-mode automatic differentiation to compute the gradients needed for parameter updating in neural networks [4].
-At the same time, training of DNNs consisting of many layers of nonlinear transformations is computationally expensive and has been associated with significant energy consumption at a global scale [5, 6, 7, 8, 9, 10].
-Hence, we are in need of green AI solutions which can significantly reduce the energy consumption of modern AI systems [11, 12, 13].
-Patterson et al.
-[14] describe several ways to reduce the carbon consumption of deep learning, focusing mostly on more effective deployment methods.
+Modern AI relies heavily on deep learning (DL), which refers to the training of very deep neural network (DNN) models using massive datasets deployed on high-performance compute clusters (LeCun et al., 2015).
+The established way of implementing learning in artificial neural networks is through the backpropagation (BP) algorithm (Linnainmaa, 1970; Werbos, 1974).
+BP is a gradient-based method that implements reverse-mode automatic differentiation to compute the gradients needed for parameter updating in neural networks (Baydin et al., 2018).
+At the same time, training of DNNs consisting of many layers of nonlinear transformations is computationally expensive and has been associated with significant energy consumption at a global scale (García-Martín et al., 2019; de Vries, 2023; Thompson et al., 2021; Luccioni et al., 2023b; Strubell et al., 2020; Luccioni et al., 2023a).
+Hence, we are in need of green AI solutions which can significantly reduce the energy consumption of modern AI systems (Gibney, 2022; Schwartz et al., 2020; Wang, 2021).
+Previous work describes several ways to reduce the carbon consumption of deep learning, focusing mostly on more effective deployment methods (Patterson et al., 2022).
 Another key reason for deep learning’s large carbon consumption, however, is the inefficiency of the backpropagation algorithm itself.
 BP requires many iterations of gradient descent steps until parameters converge to their optimal values.
-For example, training a large GPT model can easily take several weeks on a large compute cluster [15].
-In this paper, we show that deep learning can be made much more efficient by enforcing decorrelated inputs throughout the network.
-Decorrelation has previously been proposed to make credit assignment more efficient in neural networks [16].
+For example, training a large GPT model can easily take several weeks on a large compute cluster (Brown et al., 2020).
+In this paper, we show that the efficiency of deep learning can be substantially increased by enforcing that layer inputs remain decorrelated throughout the network.
+Decorrelation has previously been proposed to make credit assignment more efficient in neural networks (LeCun et al., 2012).
 Intuitively, if a network layer’s inputs have highly correlated features, it will be more difficult for the learning algorithm to perform credit assignment as it is now unclear if a change in the loss should be attributed to feature $i$ or feature $j$ in case both are correlated.
-Previous work has indeed shown that promoting decorrelation positively impacts training in relatively shallow networks [17].
-Similar results have been obtained when inputs are whitened such that inputs are not only decorrelated but also forced to have unit variance [18].
-In [19], whitening has been shown to have a positive impact in training deep networks, resulting in faster convergence when measured by number of epochs and somewhat lower test error compared to regular BP.
-However, due to computational overhead this did not yet translate into a substantial reduction in wall-clock training time.
-Desjardins et al.
-[20] also demonstrate that whitening has benefits when training DNNs, though this did not result in faster convergence to the minimal validation loss compared to standard BP using batch normalisation.
-At the same time, [21] suggest that strict whitening can have a negative impact on generalization performance.
-Hence, while the potential benefits of decorrelation and whitening are encouraging, their application for significantly reducing real-world training times in modern deep neural networks remains to be shown.
-This has two main reasons.
-First, enforcing network decorrelation during training is associated with significant computational overhead, thereby eliminating any gains in training efficiency.
-Second, it is not obvious how to effectively implement decorrelation in a stable manner in deep (convolutional) neural networks.
-In this paper we demonstrate, for the first time, that training of large-scale DNNs can be made much more efficient through the development of a novel decorrelated backpropagation (DBP) algorithm.
+Previous work has indeed shown that promoting decorrelation positively impacts training in relatively shallow networks (Ahmad et al., 2023).
+This advantage of decorrelation can be theoretically understood as a way to better align the gradient update with that of the natural gradient (Ahmad, 2024; Amari, 1998).
+Similar results have been obtained when inputs are whitened such that inputs are not only decorrelated but also forced to have unit variance (Luo, 2017).
+Whitening has also been shown to have a positive impact in training deep networks, resulting in faster convergence when measured by number of epochs (Huangi et al., 2018).
+However, since decorrelation and whitening methods themselves incur computational overhead, convergence speedup as a function of the number of epochs is not reflected in reduction in wall-clock time or carbon consumption, which is the metric of choice in green AI.
+Furthermore, other work indicates that whitening may have a convergence speed that is similar to that of batch-normalized networks (Desjardins et al., 2015) and generalization performance may be negatively impacted (Wadia et al., 2021).
+Hence, while the theoretical benefits of decorrelation and whitening are encouraging, their practical application to significantly reduce real-world training time in modern deep neural networks remains to be shown.
+In this paper we demonstrate, for the first time, that training of DNNs can be substantially improved in practice by employing a novel decorrelated backpropagation (DBP) algorithm.
 DBP combines automatic differentiation with an efficient iterative local learning rule which effectively decorrelates layer inputs across the network.
-This iterative decorrelation procedure was previously introduced by Ahmad et al.
-[17] and has been shown to speed up learning in limited-depth fully-connected neural networks.
-Here, we extend the procedure to more efficiently and effectively train deep residual networks [22].
-This is achieved by making the procedure suitable for application in convolutional layers and ensuring that decorrelation remains stable across layers.
-Furthermore, we show how the same algorithm can be used to whiten layer inputs, allowing us to compare the efficacy of decorrelation and whitening in DNNs.
-In the following, we show that DBP yields a two-fold reduction in training time, while achieving better performance compared to regular BP.
-Hence, widespread application of our approach can yield a substantial reduction in the carbon consumption of modern deep learning.
+This iterative decorrelation procedure generalizes a method which was previously introduced in the context of biologically plausible learning, demonstrating faster convergence in limited-depth fully-connected neural networks (Ahmad et al., 2023).
+Here, we extend the method to allow efficiently and effectively training of deep convolutional neural networks (He et al., 2016).
+Our method is compared on four DNN architectures trained on the ImageNet dataset.
+To ensure a fair comparison across methods, we include batch normalization and optimized hyper-parameter optimization for both methods separately.
+We show that wall-clock time of DNN training can be significantly reduced, up to 50% of that of regular BP, also reducing carbon emissions.
+Furthermore, we show that test accuracy of DBP-trained networks is consistently higher than that of BP-trained networks.
+Overall, our results show that including a local decorrelation mechanism across network layers can substantially improve training of deep neural networks.
 2 Methods
@@
 Backpropagation updates the parameters $\bm{\theta}$ (weights and biases) of a network according to
-$\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}\,,$ (1)
+$\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\bm{\theta}}\mathcal{L}$ (1)
 where $\eta$ is the learning rate, $\bm{\theta}$ are the network parameters and $\nabla_{\bm{\theta}}\mathcal{L}$ is the gradient of the loss.
-Decorrelated backpropagation (DBP) proceeds as regular backpropagation but additionally enforces decorrelated inputs to layers in the network.
+Decorrelated backpropagation (DBP) operates as regular backpropagation but additionally enforces the input to network layers to be decorrelated.
 Since decorrelation is imposed independently in each layer, we concentrate on describing the decorrelation procedure for one such layer, implementing a nonlinear transformation
 $\mathbf{y}=f\left(\mathbf{W}\mathbf{x}\right)$ (2)
-with $\mathbf{W}\in\mathbb{R}^{d\times d}$ the weights and $\mathbf{x}\in\mathbb{R}^{d}$ the input.
-This transformation may be either the kernel function of a convolutional layer applied to a (flattened) input patch or the transformation in a fully-connected layer.
-Note further that we ignore biases without loss of generality since these can be represented using weights with fixed constant input.
-Input decorrelation refers to the property that $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is diagonal for inputs $\mathbf{x}$, where the expectation is taken over the data distribution.
-In case of whitened inputs, $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is assumed to be the identity matrix [23].
+with input $\mathbf{x}\in\mathbb{R}^{d}$, weights $\mathbf{W}\in\mathbb{R}^{m\times d}$ and output $\mathbf{y}\in\mathbb{R}^{m}$.
+The nonlinear transformation may be either the transformation in a fully-connected layer or the kernel function of a convolutional layer applied to a (flattened) input patch.
+We ignore biases without loss of generality since these can be represented using weights with fixed constant input.
+Input decorrelation refers to the property that the second-moment matrix $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is diagonal for inputs $\mathbf{x}$, where the expectation is taken over the data distribution.
+In case of whitened inputs, $\langle\mathbf{x}\mathbf{x}^{\top}\rangle$ is additionally assumed to be the identity matrix (Kessy et al., 2015).
 We will refer to the former as the covariance constraint and the latter as the variance constraint.
-To ensure that the input $\mathbf{x}$ is decorrelated, we assume that it is the result of a linear transform
-$\mathbf{x}=\mathbf{R}\mathbf{z}\,,$ (3)
-where the decorrelating matrix $\mathbf{R}$ transforms the correlated input $\mathbf{z}$ into a decorrelated input $\mathbf{x}$. Hence, the transformation in a layer is described by
+To ensure that the input $\mathbf{x}$ is decorrelated, we assume that it is the result of a linear transform $\mathbf{x}=\mathbf{R}\mathbf{z}$, where the decorrelating matrix $\mathbf{R}$ transforms the correlated input $\mathbf{z}$ into a decorrelated input $\mathbf{x}$.
+Hence, the transformation in a layer is described by
 $\mathbf{y}=f(\mathbf{W}\mathbf{R}\mathbf{z})=f(\mathbf{W}\mathbf{x})=f(\mathbf{A}\mathbf{z})$ (3)
@@
 To implement decorrelated backpropagation, we need to update individual decorrelation matrices for each of the layers in the network.
-This can be achieved by minimizing $K$ decorrelation loss functions in parallel to minimizing the BP loss with the aim of accelerating learning speed.
-One way to achieve whitening as a restricted form of decorrelation is via established approaches such as zero-phase component analysis (ZCA) [24, 25].
-However, naive application of ZCA is prohibitively costly in practice since it requires application of the algorithm to the inputs of each layer for all of the examples in the dataset at each step of gradient descent.
-More efficient batch implementations of whitening have been shown to have a positive impact on neural network training [18, 20, 19] but this did not yet yield a substantial reduction in training time in large-scale models when considering the moment at which minimal validation loss is achieved.
-See Appendix A for a complexity analysis of the different approaches.
-To enable efficient decorrelation and whitening, we extend a recently introduced approach that enables more efficient network-wide decorrelation [17] and make this suitable for application in large-scale models.
-To derive our decorrelation learning rule, we start by defining the total decorrelation loss $\langle L\rangle$ as the average decorrelation loss across layers.
+This can be achieved by minimizing $K$ decorrelation loss functions in parallel with minimizing the BP loss with the aim of accelerating learning speed.
+One way to achieve whitening as a restricted form of decorrelation is via established approaches such as zero-phase component analysis (ZCA) (Bell and Sejnowski, 1995, 1997).
+However, as described in Appendix A, naive application of ZCA is prohibitively costly in practice since it requires computing an expensive ZCA transform for the inputs of each layer at each gradient-descent step.
+To enable efficient decorrelation and whitening in DNNs, we generalize a recently introduced approach that allows for efficient network-wide decorrelation (Ahmad et al., 2023).
+To derive the decorrelation learning rule, we start by defining the total decorrelation loss as the average decorrelation loss across layers.
 The layer-specific decorrelation loss is defined as $L=\langle\ell(\mathbf{x})\rangle$, where the expectation is taken over the empirical distribution and $\ell(\mathbf{x})=\sum_{i=1}^{d}\ell_{i}(\mathbf{x})$ is the sum over unit-wise losses $\ell_{i}$.
 We define the unit-wise loss as
-$\ell_{i}(\mathbf{x})=(1-\kappa)\frac{1}{2}\sum_{j\neq i}(x_{i}x_{j})^{2}+\kappa\frac{1}{4}(x_{i}^{2}-1)^{2}\,,$ (5)
-where the regularization parameter $\kappa$ interpolates between the covariance constraint ($\kappa=0$) and the unit variance constraint ($\kappa=1$) such that that $\kappa\in(0,1)$ promotes whitening.
-It follows that we may write the layer-wise decorrelation loss compactly as
-$\ell(x)=\left\langle(1-\kappa)\textrm{Tr}(\mathbf{C}\mathbf{C}^{\top})+\kappa\textrm{Tr}(\mathbf{V}\mathbf{V}^{\top})\right\rangle\,,$ (6)
+$\ell_{i}(\mathbf{x})=(1-\kappa)\frac{1}{2}\sum_{j\neq i}(x_{i}x_{j})^{2}+\kappa\frac{1}{4}(x_{i}^{2}-1)^{2}$ (4)
+where the first term is the zero covariance constraint imposed by decorrelation and the second term is the unit variance constraint imposed by whitening.
+The whitening parameter $\kappa\in[0,1]$ trades off between the covariance constraint ($\kappa=0$) and the variance constraint ($\kappa=1$) and is whitening inducing when $\kappa>0$.
+we may now write the layer-wise decorrelation loss compactly as
+$\ell(\mathbf{x})=\left\langle(1-\kappa)\textrm{Tr}(\mathbf{C}\mathbf{C}^{\top})+\kappa\textrm{Tr}(\mathbf{V}\mathbf{V}^{\top})\right\rangle$ (5)
 where $\mathbf{C}=\mathbf{x}\mathbf{x}^{\top}-\mathbf{D}$ with $\mathbf{D}=\textrm{diag}(x_{1}^{2},\ldots,x_{d}^{2})$ and $\mathbf{V}=\textrm{diag}(x_{1}^{2}-1,\ldots,x_{d}^{2}-1)$.
@@
 with $\epsilon$ the decorrelation learning rate. To compute the gradient, we first consider the partial derivative of the unit-wise loss with respect to the input $x_{i}$ given by
-$\frac{\partial\ell_{i}}{\partial x_{i}}=(1-\kappa)\sum_{j\colon j\neq i}\left(x_{i}x_{j}\right)x_{j}+\kappa(x_{i}^{2}-1)x_{i}\,.$ (8)
-If we vectorize across units, we obtain $\nabla_{\mathbf{x}}\ell=(1-\kappa)\mathbf{C}\mathbf{x}+\kappa\mathbf{V}\mathbf{x}$. If we plug this into the update step in Eq. 7, we obtain
-$\mathbf{x}\leftarrow\mathbf{x}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{x}\,.$ (9)
-Ultimately, we aim to derive an update rule for $\mathbf{R}$ instead of $\mathbf{x}$.
+$\frac{\partial\ell_{i}}{\partial x_{i}}=(1-\kappa)\sum_{j\colon j\neq i}\left(x_{i}x_{j}\right)x_{j}+\kappa(x_{i}^{2}-1)x_{i}$ (7)
+which, when vectorized across inputs, yields $\nabla_{\mathbf{x}}\ell=(1-\kappa)\mathbf{C}\mathbf{x}+\kappa\mathbf{V}\mathbf{x}$.
+If we now plug this into the update step in Eq.
+6, we obtain $\mathbf{x}\leftarrow\mathbf{x}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{x}$.
+Note, however, that our goal is to derive an update rule for $\mathbf{R}$ rather than $\mathbf{x}$.
 This can be obtained using the identity $\mathbf{x}=\mathbf{R}\mathbf{z}$ since this allows us to write $\mathbf{R}\mathbf{z}\leftarrow\mathbf{R}\mathbf{z}-\epsilon\left((1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right)\mathbf{R}\mathbf{z}$.
-Dividing by $\mathbf{z}$ on both sides and averaging over input samples we obtain our decorrelation learning rule
+Dividing by $\mathbf{z}$ on both sides and averaging over input samples we obtain the decorrelation learning rule
 $\mathbf{R}\leftarrow\mathbf{R}-\epsilon\left\langle(1-\kappa)\mathbf{C}+\kappa\mathbf{V}\right\rangle\mathbf{R}\,.$ (8)
 This decorrelation rule allows for efficient batch updating of $\mathbf{R}$ since it only requires computing the decorrelated input covariance followed by multiplication with the decorrelation matrix.
-Note that for $\kappa=0$ we only decorrelate the input and for $\kappa=1$ we only enforce the unit variance constraint.
-For $10<\kappa<$ 1 we include both constraints to induce input whitening.
-Figure 1 shows effective decorrelation and whitening on two-dimensional input data.
+Figure 1 illustrates effective decorrelation and whitening on two-dimensional input data using Eq. (8).
 Figure 1: Demonstration of the decorrelation rule on correlated input data consisting of 1000 examples and two covariates with decorrelation learning rate $\epsilon=0.001$.
@@
 Mean variance and covariance reported for different iterations.
-2.3 Application to deep residual networks
-Effective training of deep neural networks using DBP requires several modifications that improve algorithm stability, learning speed and allow application to convolutional rather than fully connected layers.
-As a canonical example, we consider an 18-layer deep residual network (ResNet).
-A ResNet consists of multiple residual blocks that implement modern network components such as convolutional layers and residual connections [22].
+2.3 Decorrelating deep convolutional neural networks
+Effective training of DNNs using DBP requires several modifications that improve algorithm stability, learning speed and allow application to convolutional rather than fully connected layers.
+As a canonical example, we consider a deep residual network (ResNet).
+A ResNet consists of multiple residual blocks that implement modern network components such as convolutional layers and residual connections (He et al., 2016).
 In the following, we describe the modifications that are required for effectively applying DBP in deep neural networks.
@@
 a) Residual blocks as implemented by our networks.
-b) Patch-wise flattening of the inputs with a flattened dimension of $d=M\times M\times C_{in}$.
+b) Patch-wise flattening of the inputs with a flattened dimension of $d=M\times M\times C_{\text{in}}$.
 c) Decorrelating/whitening transform of the data by decorrelation matrix $\mathbf{R}$.
 d) 1 $\times$ 1 convolution operation with the weights $\mathbf{W}$ on the decorrelated input patches.
-First, we apply the decorrelation learning rule to convolutional layers as follows.
-Consider a convolutional layer with input size $S=M_{in}\times M_{in}\times C_{in}$, where $M_{in}\times M_{in}$ is the size of the feature map and $C_{in}$ is the channel dimension.
-Naive implementation of decorrelation would require updating a $S\times S$ decorrelation matrix $\mathbf{R}$, which is computationally too prohibitive in case of large feature maps.
+Decorrelating convolutional layers
+We apply the decorrelation learning rule to convolutional layers as follows.
+Consider a convolutional layer with input size $S=M_{\text{in}}\times M_{\text{in}}\times C_{\text{in}}$, where $M_{\text{in}}\times M_{\text{in}}$ is the size of the feature map and $C_{\text{in}}$ is the channel dimension.
+Naive implementation of decorrelation would require updating a $S\times S$ decorrelation matrix $\mathbf{R}$, which is computationally prohibitive in case of large feature maps.
 Instead, as shown in Fig.
-2, we do not decorrelate the layer’s entire input, but only the local image patches, so that for an image patch with dimensionality $D=K\times K\times C_{in}$, matrix $\mathbf{R}$ would be of much smaller size $D\times D$.
-The output of this patchwise decorrelation operation would then be $D$ for every image patch, after which we can apply a $1\times 1$ convolutional kernel of size $D\times C_{out}$ for the forward pass.
+2, we do not decorrelate the layer’s entire input, but only the local image patches, so that for an image patch with dimensionality $D=K\times K\times C_{\text{in}}$, matrix $\mathbf{R}$ would be of much smaller size $D\times D$.
+The output of this patchwise decorrelation operation would then be $D$ for every image patch, after which we can apply a $1\times 1$ convolutional kernel of size $D\times C_{\text{out}}$ for the forward pass.
 This local decorrelation is sufficient to ensure more efficient learning of the kernel weights.
 Figure 2 depicts the structure of a residual block that is extended to implement decorrelating transforms.
+Subsampling during covariance estimation
 Second, applying the decorrelation learning rule can become costly when averaging over input samples in Eq.
@@
 We found empirically that updates of $\mathbf{R}$ need not use the entire mini-batch to learn the correlational structure of the data.
-Sampling just 10% of the samples in each batch yields almost identical performance while significantly reducing the computational overhead of computing $\mathbf{R}$’s updates.
-Third, another optimization that decreases the computational overhead of the decorrelating transform is to combine the matrices $\mathbf{W}$ and $\mathbf{R}$ into a condensed matrix $\mathbf{A}=\mathbf{W}\mathbf{R}$ prior to multiplying by the correlated input $\mathbf{z}$.
+Sampling just 10% of the samples in each batch yields almost identical performance while significantly reducing the computational overhead of computing $\mathbf{R}$’s updates, as shown in Appendix B.
+Efficient matrix products
+A final optimization that decreases the computational overhead of the decorrelating transform is to combine the matrices $\mathbf{W}$ and $\mathbf{R}$ into a condensed matrix $\mathbf{A}=\mathbf{W}\mathbf{R}$ prior to multiplying by the correlated input $\mathbf{z}$.
 The dimensionality of $\mathbf{z}$ is much larger than that of $\mathbf{W}$ due to the batch dimension, and thus we replace an expensive multiplication $\mathbf{W}(\mathbf{R}\mathbf{z})$ by a cheaper multiplication $(\mathbf{W}\mathbf{R})\mathbf{z}$, significantly reducing the time required for performing a forward pass.
@@
 2.4 Experimental validation
-To evaluate learning performance, an 18-layer deep ResNet model was trained and tested on the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) dataset [26].
-This dataset spans 1000 object classes and contains 1,281,167 training images, 50,000 validation images and 100,000 test images [26].
+To evaluate learning performance in DNNs, 18-, 34- and 50-layer ResNet models as well as an AlexNet model (Krizhevsky et al., 2012) were trained using both BP and DBP.
+Performance was evaluated using the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) dataset (Russakovsky et al., 2015).
+This dataset spans 1000 object classes and contains 1,281,167 training images, 50,000 validation images and 100,000 test images.
 Data preprocessing consisted of the following steps.
@@
 We used no data augmentation, but shuffled the order of the data every epoch.
-All models and algorithms were implemented in PyTorch [27] and run on a compute cluster using Nvidia A100 GPUs and Intel Xeon Platinum 8360Y processors.
-To initialize the weights of our models, we set $\mathbf{R}$ to the identity matrix and used He initialization [28] for $\mathbf{W}$.
-Models were trained for 50 epochs to minimize the categorical cross-entropy loss.
+All models and algorithms were implemented in PyTorch (Paszke et al., 2019) and run on a compute cluster (Surf Snellius) using Nvidia A100 GPUs and Intel Xeon Platinum 8360Y processors.
+To initialize the weights of our models, we set $\mathbf{R}$ to the identity matrix and used He initialization (He et al., 2015) for $\mathbf{W}$.
+Models were trained to minimize the categorical cross-entropy loss.
 Exploratory analysis revealed negligible performance difference as a function of batch size and a batch size of 256 was chosen.
@@
 We empirically found this value to show negligible loss in performance, while significantly speeding up runtime.
+See Appendix B for an analysis.
 The accuracy reported is the top-1 performance of the models.
 Reported wall-clock time was measured as the runtime of the training process, excluding performance testing.
-Performance of DBP using either decorrelation ($\kappa=0$) or whitening ($\kappa=0.5$) was compared against regular backpropagation.
-For updating $\mathbf{W}$, we use the Adam optimizer [29] with a learning rate of $\eta=1.6\cdot 10^{-4}$ and optimizer parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
+Total compute time used is estimated at 400 hours, which includes initial exploration, hyper-parameter search and running the large-scale experiments.
+For the $\kappa$ parameter, a value of 0.5 was chosen, balancing the covariance and variance constraints.
+For updating $\mathbf{W}$, we use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $\eta=1.6\cdot 10^{-4}$ and optimizer parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
 We added a small value of $10^{-8}$ to the denominator of the updates for numerical stability.
@@
 To ensure a fair comparison between algorithms, we made sure that all algorithms attained their highest convergence speed by performing a two-dimensional grid search over the BP ($\eta$) and decorrelation ($\epsilon$) learning rates.
-See Section B for the grid-search results.
+See Appendix C for the grid search results on the ResNet18 model.
 3 Results
-In the following, we analyse the performance of BP versus DBP on the 18-layer ResNet model trained on the ImageNet task.
+In the following, we analyze the performance of DBP on different DNN models trained on the ImageNet task.
 3.1 DBP effectively decorrelates inputs to all network layers
-In Fig.
-3, the change is input correlation is shown for each layer during the training process for DBP.
+We first verified that DBP indeed effectively reduces input decorrelation across layers in DNNs.
+Figure 3 shows the change is input correlation for each layer of the ResNet18 model during the training process for DBP.
 To quantify input correlation we use the mean squared values of the strictly lower triangular part of the $\mathbf{C}$ matrix.
-It can be seen that layer inputs become rapidly decorrelated during the first 10 epochs and correlations remain consistently low.
+Layer inputs become rapidly decorrelated during the first 10 epochs and correlations remain consistently low.
 This demonstrates that the decorrelating learning rule indeed effectively learns to decorrelate the inputs to all network layers.
-Figure 3: Input decorrelation when training a ResNet-18 model on ImageNet for 50 epochs. Network layers are ordered from left to right and from top to bottom. Panel titles indicate the layer type.
+Figure 3: Input decorrelation when training a ResNet18 model on ImageNet for 50 epochs. Network layers are ordered from left to right and from top to bottom. Panel titles indicate the layer type.
 3.2 DBP converges much faster than BP
-We observe a significant improvement when training the ResNet model using DBP compared to regular BP.
-As shown in Fig.
-4, DBP’s train accuracy (Fig.
-4a) improves much faster than that of BP, indicating more effective training (computational work per epoch).
-The same observation holds for test accuracy (Fig.
-4b).
-Additionally, DBP’s test accuracy (55.2%) also peaks above BP’s test accuracy (54.1%), while doing so in roughly half the number of epochs.
-Finally, we observe that a slightly higher test accuracy is achieved for whitening ($\kappa=0.5$) compared to decorrelation ($\kappa=0$), demonstrating the added benefit of whitening for this task.
-The same patterns can be observed for the loss curves in Appendix C.
-Furthermore, as shown in Appendix D, a slightly lower test loss at equal train loss is obtained for DBP compared to BP, which shows that this does not come at the expense of generalization performance.
-Figure 4: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening).
+Figure 4 compares training and test convergence using DBP and BP as a function of the number of epochs across models.
+For each model, convergence of both train and test accuracy is much faster for DBP compared to BP, demonstrating the positive impact of decorrelation.
+Furthermore, test performance is consistently higher for DBP compared to BP.
+Figure 4: Train and test performance of BP and DBP on Imagenet as a function of epochs for different deep neural network architectures.
+AlexNet was trained for 20 epochs and the deeper ResNet models were trained for 35 epochs.
+a) AlexNet.
+b) ResNet18.
+c) ResNet34.
+d) ResNet50.
+3.3 DBP training yields shorter wall-clock times
+As emphasized in the Introduction section, speedup in terms of epochs may not necessarily translate into speedup in terms of wall-clock time due to additional computational overhead.
+Figure 5 shows that, while the speedup is indeed attenuated due to this overhead, DBP remains faster than BP, reducing training time up to 50% across models when using peak test performance as a stopping criterion.
+Figure 5: Train and test performance of BP and DBP on Imagenet as a function of wall-clock time for different deep neural network architectures.
+AlexNet was trained for 20 epochs and the deeper ResNet models were trained for 35 epochs.
+a) AlexNet.
+b) ResNet18.
+c) ResNet34.
+d) ResNet50.
+3.4 DBP reduces carbon emission
+Reduced wall-clock time does not only reduce the need for compute capacity but also decreases the carbon emission of deep learning.
+Figure 6 summarizes our results, showing faster learning at higher accuracy with reduced carbon consumption.
+Figure 6: a) Maximum test accuracy b) Time needed to achieve maximum test accuracy c) Estimated reduction in carbon consumption when using DBP over BP.
+Estimates produced using https://github.com/mlco2/codecarbon.
+3.5 Impact of whitening
+The question remains how the choice of the whitening parameter $\kappa$ influences performance.
+To this end, we compared the performance of the ResNet18 model using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening) for three random initializations.
+Figure 7 show that performance is slightly improved for $\kappa=0.5$ compared to $\kappa=0$.
+However, as shown in Appendix D, the optimal choice of $\kappa$ depends on the chosen architecture and dataset.
+Figure 7: Performance of BP and DBP on ResNet18 using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening).
 Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars.
@@
 b) Test accuracy as a function of the number of epochs.
-3.3 DBP training can be achieved at shorter wall-clock times than BP
-Recall that DBP requires additional computation in each epoch due to application of the decorrelation update rule in Eq.
-10.
-This implies that faster convergence as a function of the number of epochs may not necessarily translate into more efficient training.
-Therefore, we also compare wall-clock times in Fig.
-5.
-Figure 5: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening).
-Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars.
-a) Train accuracy as a function of wall-clock time.
-b) Test accuracy as a function of wall-clock time.
-[…]
-3.4 DBP significantly reduces carbon emission
-Figure 6 visualizes what the efficiency gains are when using DBP ($\kappa=0.5$) over BP in terms of carbon emission.
-Figure 6: Efficiency gains when using DBP over BP.
-a) Training time required relative to BP’s time to maximal accuracy to reach a specific percentage of the maximum performance of DBP.
-b) Estimated reduction in carbon consumption when using DBP over BP.
-Estimate produced using https://github.com/mlco2/codecarbon.
-Dashed line indicates the moment when performance is equal to maximal BP test accuracy.
-[…]
-Figure 6b provides a measurement of the associated reduction in carbon emission if we run the model to achieve a specific percentage of the DBP test accuracy.
-The figure shows that a reduction of 640 grams of carbon emission is achieved if we train DBP until it reaches 100% of the BP test accuracy.
-If we run DBP until it reaches 100% of the DBP test accuracy, we still achieve a reduction of 418 grams of carbon emission.
 4 Discussion
 Increasing the efficiency of deep learning is of key importance if we are to reduce the carbon footprint of AI.
-In this paper, we have shown that decorrelated backpropagation provides a viable path towards more efficient deep learning.
-Results show that, by replacing BP with DBP, we can achieve higher test accuracy using about half the number of training epochs.
-Results also show that both decorrelation ($\kappa=0$) and whitening ($\kappa=0.5$) provide similar performance gains, with whitening yielding slightly better results.
-This does not come at the expense of generalization performance, as shown in Fig.
-9.
-We also demonstrated a substantial reduction in carbon emission when using DBP over BP when training deep neural networks.
-Global carbon emissions can therefore be substantially reduced when applying DBP at scale.
-These results indicate that performance is maximized for some optimal value of $\kappa$ which most effectively balances the covariance and variance constraints.
-This optimal value is task-dependent.
-Appendix F shows that training a medium-sized neural network on the CIFAR10 task yields similar performance gains but now the optimal performance is obtained for $\kappa=0$.
-It should be mentioned that network-wide decorrelation comes with a number of additional advantages, which we did not further explore here.
-As shown in [17], decorrelated inputs allow for rapid linear approximation of non-linear functions, allowing for filter visualisation in deep networks with applications in explainable AI [30, 31].
-It also allows for network compression at virtually no computational overhead, which may further decrease carbon consumption at inference time [13].
-Intriguingly, input decorrelation and whitening have also been shown to be a feature of neural processing [32, 25, 33, 34, 35, 36, 37].
-The work presented here suggests that decorrelation may be an important factor when considering synaptic plasticity mechanisms.
-A number of considerations do need to be taken into account when using DBP.
-First, DBP requires updating and storing of decorrelation matrices needed for input decorrelation.
-Hence, the algorithm introduces some computational overhead and increased memory requirements.
-This is reflected by the increased wall-clock time per epoch.
-To mitigate this, more effective use could be made of the sparseness structure of the decorrelation matrix.
+In this paper, using four deep neural network architectures trained on the Imagenet dataset, we have shown that decorrelated backpropagation, which decorrelates the inputs to each network layer, provides a viable path towards more efficient deep learning.
+Results show that, by replacing BP with DBP, we achieve higher test accuracy using up to half the number of training epochs, resulting in reduced carbon emission.
+The observed reduction in training time and associated carbon emission of DBP is the result of a complex interplay between having a closer alignment of the gradient with the natural gradient, yielding faster convergence per epoch, and the need to decorrelate inputs across network layers, which comes with additional overhead per epoch.
+This overhead stems from a number of factors which apply to all $1\leq k\leq K$ network layers.
+First, the need to store the decorrelation matrices $\mathbf{R}_{k}$.
+Second, computing the gradient update (8) for each matrix $\mathbf{R}_{k}$ in parallel.
+Third, additional matrix multiplications $\mathbf{R}_{k}$ and $\mathbf{R}^{\top}_{k}$ during the forward and backward pass of the backpropagation algorithm.
+These factors contribute to increased memory and compute requirements though this is more than compensated for by the much faster convergence of DBP compared to BP.
+These considerations suggest that even larger performance gains may be achievable if we can further reduce the memory and compute requirements of DBP.
+One way to achieve this may be to exploit sparseness structure in the decorrelation matrix.
 For instance, rather than learning a full matrix $\mathbf{R}$ we may choose to learn a lower-triangular matrix, which can be achieved by using the strictly lower triangular part of $\mathbf{C}$ in Eq.
-10.
-Other approaches to reduce wall-clock time may be the use of low-rank approximations of $\mathbf{R}$, using higher layer-specific decorrelation learning rates, or only incorporating decorrelation in that subset of the layers which have the biggest impact in terms of reducing convergence time [19].
-Second, DBP needs proper fine-tuning since the stability of the algorithm depends on the two learning rates $\eta$ and $\epsilon$, whose ratio needs to be set appropriately such that the decorrelation rate throughout all network layers remains optimal.
+8.
+Another approach may be to use a low-rank approximation of $\mathbf{R}$.
+Finally, we may choose to only incorporate decorrelation in that subset of layers which have the biggest impact in terms of reducing convergence time (Huangi et al., 2018).
+Further speedups may be achievable by improving the decorrelation algorithm itself, for instance, by using layer-specific decorrelation learning rates or subsampling rates that take layer properties into account.
+We expect that additional theoretical work on the optimal alignment of layer inputs combined with efficient low-level implementations will yield even greater gains in convergence speed.
+It is important to realize that DBP needs proper fine-tuning.
+The stability of the algorithm depends on the interplay between the backpropagation learning rate $\eta$ and the decorrelation learning rate $\epsilon$.
+This ratio needs to be set appropriately such that the decorrelation rate throughout all network layers remains optimal.
+That is, a too slow decorrelation learning rate will not improve convergence speed while still incurring computational overhead.
+On the other hand, a too fast decorrelation learning rate will reduce the ability of backpropagation to keep up with changes in input statistics across the network, leading to instabilities.
 We empirically observe that this fine-tuning depends on the employed dataset, network architecture and loss function.
 Additional work on the interaction between these factors may provide insights into how to optimally set these parameters.
-An important consideration here is scaling the updates according to layer size.
+Our results further show that performance is maximized for some optimal value of $\kappa$ which most effectively balances the covariance and variance constraints.
+As a rule of thumb, the setting $\kappa=0.5$ is reasonable though we may also choose to set $\kappa=0$, further simplifying the decorrelation learning rule to $\mathbf{R}\leftarrow\mathbf{R}-\epsilon\left\langle\mathbf{C}\right\rangle\mathbf{R}$.
+Another important consideration is normalizing the decorrelation update magnitude according to layer size.
 Note that Eq.
-10 ignores the layer size in the normalisation, which implies that decorrelation updates have a larger impact in larger layers whereas whitening induced by $\kappa>0$ has a smaller impact in larger layers.
-In simulation work, we observe that such normalisation ensures that the impact of decorrelation updates is comparable for different layer sizes.
-However, in our empirical work using very deep neural networks, we find that ignoring this normalisation actually improves performance.
+8 ignores the layer size in the normalization, which implies that decorrelation updates have a larger impact in larger layers whereas whitening induced by $\kappa>0$ has a smaller impact in larger layers.
+In simulation work, we observe that such normalization ensures that the impact of decorrelation updates is comparable for different layer sizes.
+However, in our empirical work using very deep neural networks, we find that disregarding this normalization actually improves performance.
 There are different reasons why this may be the case.
-First, the resulting stronger decorrelation in deep large layers when ignoring normalisation may have a beneficial impact since deeper layers are more strongly affected by multiple preceding nonlinearities.
+First, the resulting stronger decorrelation in deep large layers when ignoring normalization may have a beneficial impact since deeper layers are more strongly affected by multiple preceding nonlinearities.
 Second, the decorrelation loss that is minimized by Eq.
 8 may not be the optimal metric to ensure fastest convergence.
-That is, there may be a discrepancy between the decorrelation gradient direction and the optimal gradient direction to align the inputs.
-In this work, we demonstrated a significant speedup in deep residual networks, which are state-of-the-art models for computer vision.
-In our follow-up work, we aim to explore the efficiency gains that can be obtained for other computational tasks and network architectures.
-Exciting prospects are the use of decorrelated backpropagation for deep reinforcement learning [38, 39] or training of foundation models [40], both of which are notoriously resource intensive.
-We further expect that additional theoretical work on the optimal alignment of layer inputs combined with efficient low-level implementations will yield even larger gains in convergence speed.
+That is, there may be an as yet unidentified discrepancy between the decorrelation gradient direction and the optimal gradient direction to align the inputs.
+It should be mentioned that network-wide decorrelation comes with a number of additional advantages, which we did not further explore here.
+As shown in (Ahmad et al., 2023), decorrelated layer inputs allow for extremely efficient computation of a optimal linear mapping between arbitrary network layers.
+That is, when inputs are decorrelated, the matrix that needs to be inverted to compute that ordinary least squares solution becomes diagonal, reducing the inversion from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d)$ with $d$ the number of inputs.
+This property allows for filter visualization in deep networks with applications in explainable AI (Rudin, 2019; Ras et al., 2022) as well as for network compression at virtually no computational overhead, which may further decrease carbon consumption at inference time (Wang, 2021).
+Intriguingly, input decorrelation and whitening have also been shown to be a feature of neural processing (Franke et al., 2017; Bell and Sejnowski, 1997; Pitkow and Meister, 2012; Segal et al., 2015; Dodds et al., 2019; Graham et al., 2006; King et al., 2013).
+The work presented here suggests that decorrelation may be an important factor when considering synaptic plasticity mechanisms, warranting further investigation.
+In this work, we demonstrated a significant speedup in deep feedforward neural networks, which are state-of-the-art models for computer vision.
+In follow-up work, we aim to explore the efficiency gains that can be obtained for other computational tasks and network architectures.
+Exciting prospects are the use of decorrelated backpropagation for deep reinforcement learning (Mnih et al., 2015; Schrittwieser et al., 2020) or training of generative models (Vaswani et al., 2017; Gu and Dao, 2024; Bommasani et al., 2021), both of which are notoriously resource intensive.
 Concluding, by demonstrating the efficiency of decorrelated backpropagation in modern deep neural networks, we hope to contribute to reducing the energy consumption of modern AI systems.
-Code availability
-Code for training and evaluating the decorrelated backpropagation algorithm can be downloaded from https://github.com/artcogsys/decorbp.
 Acknowledgements
@@
 Appendix A Comparison with zero-phase component analysis
-Zero-phase component analysis (ZCA) [24, 25] is defined by the whitening transform $\mathbf{x}=\mathbf{C}\mathbf{x}_{c}$, where $\mathbf{x}_{c}=\bar{\mathbf{x}}-\bm{\mu}$ with $\bm{\mu}$ the mean over datapoints and $\mathbf{C}$ is the whitening transform.
+Zero-phase component analysis (ZCA) [Bell and Sejnowski, 1995, 1997] is defined by the whitening transform $\mathbf{x}=\mathbf{C}\mathbf{x}_{c}$, where $\mathbf{x}_{c}=\bar{\mathbf{x}}-\bm{\mu}$ with $\bm{\mu}$ the mean over datapoints and $\mathbf{C}$ is the whitening transform.
 Let $\mathbf{X}$ be the $n\times d$ matrix consisting of $n$ $d$-dimensional centered inputs.
@@
 This provides the benefit of not requiring any matrix inversion computation and also ensures that we are continuously optimising the decorrelation transform rather than doing so at fixed intervals.
-Appendix B Hyper-parameter optimization
+Appendix B Impact of subsampling
+Figure 8 shows test performance as a function of wall-clock time for AlexNet trained on ImageNet at different sampling frequencies and sampling rates.
+Sampling 10% of the batch appears sufficient for robust decorrelation while keeping computational overhead limited.
+Trying to save even more computation time by only updating the decorrelation matrix every 5 or 10 batches does not appear to speed up convergence in this experiment.
+Based on these results, a sampling rate of 10% was used in all experiments, while the sampling frequency was not adjusted.
+Figure 8: AlexNet test accuracy on ImageNet a) Sampling every batch, 5 batches, and 10 batches. b) Sampling 100%, 50%, and 10% of the batch.
+Appendix C Hyper-parameter optimization
 A grid search across learning rate hyper-parameters was performed to ensure that each algorithm reached its optimal performance.
@@
 Therefore, the latter was chosen for all algorithms, which had the convenient side effect of using the same learning rate for all algorithms, making comparison more straightforward.
-Figure 7 depicts the grid search outcomes.
+Figure 9 depicts the grid search results for the ResNet18 model.
 Figure 9: Train accuracy after five epochs for DBP when varying the BP learning rate $\eta$ and decorrelation learning rate $\epsilon$.
@@
 b) Results using whitening ($\kappa=0.5$).
-Appendix C Loss curves
-Figure 8 reports the loss curves as a function of epochs and wall-clock time for the different algorithms, corresponding to the accuracies reported in the main text.
-Figure 8: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening).
-Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars.
-a) Train loss as a function of the number of epochs.
-b) Test loss as a function of the number of epochs.
-c) Train loss as a function of wall-clock time.
-d) Test loss as a function of wall-clock time.
-Appendix D Train versus test loss comparison
-Figure 9 shows the train versus test losses for a ResNet-18 model trained on ImageNet for both BP and DBP.
-As DBP’s test loss is slightly lower than the BP test loss at the same value of the train loss, we conclude that DBP yields slightly better generalization performance compared to BP.
-Figure 9: Scatterplot of train versus test loss for BP and DBP.
-Appendix E Peak performance comparison
-Table 1 provides an overview of peak performance for the different algorithms, as well as the epoch number and time at which peak test performance was achieved.
-Table 1: Peak performance for BP, DBP in terms of loss and accuracy.
-Epoch and Time indicate the epoch number and wall-clock time at which peak accuracy was achieved.
-Best results shown in boldface, where we focus on test results as this quantifies generalization performance.
-Loss Accuracy Epoch Time (hours)
-Method $\kappa$ Train Test Train Test Train Test Train Test
-BP - 1.24 2.90 99.4 54.1 49 12 16.0 3.7
-DBP 0 1.21 2.86 99.5 54.8 50 7 18.4 2.3
-DBP 0.5 1.22 2.84 99.5 55.2 50 7 18.4 2.3
-Appendix F Impact of $\kappa$ parameter on CIFAR10 performance
-Figure 10 shows train and test performance for a small three-layer ConvNet trained on CIFAR10 for BP and for several settings of DBP where the $\kappa$ parameter was varied.
-Setting $\kappa=0$ reached the highest peak test accuracy for this particular task.
-Performance gains for this smaller model are even more pronounced yet result in a smaller reduction in carbon consumption due to the model’s smaller size.
-Note further that peak test performance for DBP is larger than that of BP, again indicating improved generalization performance.
-A batch size of 256 and learning rate of $1\times 10^{-3}$ were used in all experiments.
-The decorrelation/whitening learning rate was also set to $1\times 10^{-3}$.
+Appendix D Impact of whitening for ConvNets on CIFAR-10
+To investigate the effect of the $\kappa$ parameter in a different setting than large (residual) DNNs, a small three-layer ConvNet was trained on the CIFAR-10 [Krizhevsky, 2009] dataset.
+The model was trained using BP as welll as DBP under different values of the $\kappa$ parameter.
+As shown in Figure 10, a setting of $\kappa=0$ reached the highest peak test accuracy, though differences were small for for values less than $\kappa=0.6$.
+Note further that performance gains for this smaller model are even more pronounced than for the larger models.
 Figure 10: Performance on CIFAR10 for BP and several $\kappa$ settings of DBP. a) Train accuracy. b) Test accuracy.

### v5

@@
 To achieve this goal we made use of a novel algorithm which induces network-wide input decorrelation using minimal computational overhead.
-By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training several deep networks up to a 50-layer ResNet model.
+By combining this algorithm with careful optimizations, we achieve a more than two-fold speed-up and higher test accuracy compared to backpropagation when training several deep residual networks.
 This demonstrates that decorrelation provides exciting prospects for efficient deep learning at scale.
@@
 b) Test accuracy as a function of the number of epochs.
+3.6 Robustness to data regimes
+The results in Sections 3.2 and 3.3 intentionally show performance in a setup without data augmentation or other methods to aid performance, as the aim was to make a clean comparison of the convergence properties of the algorithms.
+In this section, we will test the robustness of our findings by exposing both BP and DBP to two additional data regimes: data augmentation consisting of random cropping and horizontal flipping as described in He et al.
+(2016), as well as downsampling the training set to 10%, stratified over classes.
+Figure 8: Test performance of BP and DBP ($\kappa=0.5$) on ResNet18 using for the regular ImageNet dataset (solid line), data augmentation (dashed line) and data downsampled to 10% (dotted line).
+Results in Figure 8 indicate that DBP’s outperformance of BP is robust in all three data regimes.
+Data augmentation improved test performance over baseline for both algorithms, which now reach 65% top-1 test accuracy, close to published benchmarks He et al.
+(2016).
+Data downsampling obviously hurt test performance for both algorithms, but DBP’s higher maximum test accuracy and faster convergence remain intact.
 4 Discussion
@@
 Results show that, by replacing BP with DBP, we achieve higher test accuracy using up to half the number of training epochs, resulting in reduced carbon emission.
+Results appear to be robust under both data augmentation and data downsampling.
 The observed reduction in training time and associated carbon emission of DBP is the result of a complex interplay between having a closer alignment of the gradient with the natural gradient, yielding faster convergence per epoch, and the need to decorrelate inputs across network layers, which comes with additional overhead per epoch.